IN a traingle A,B AND C.then proove TANA +TANB+TANC=?
 

Grade:12

12 Answers

SAI SARDAR
8 years ago
Gowri Sankar, In triangle abc a+b+c=180 then tan a + tan b + tanc = tanatanbtanc.It is from trigonametric identity.
Prabhakar ch
8 years ago
Dear Gowri Shankar
                                It would prove by some trignometric idinaties
This image tells how to proove that formulae
All the best for your bright future
N JYOTHEESWAR
8 years ago
dear gowrisankar
                        the total angle in the triangle is 180.in triangle a+b+c=180. then the angles tana+tanb+tanc=180.ie .. tanatanbtanc.this is from trigonometric identities.
T.kumar
8 years ago
dear gowri shankar ,
                                  in this triangle a b c total angels is 180 indiviusally one is 60. 2nd is 60. 3rd is 60. all the best.
Gaddam Chethan
8 years ago
Hello Gowri
                      we know that a+b+c=180. with these tanA+tanB+thaC=tanAtanBtanC. Aee the best for your bright future.
raj
8 years ago
hellow gowrishankar if A,B and C in a triangle then tanA+tanB+tanC=tanA .tanB.tanC this is proves through angles of trigonametri.
NAGOORVALI
8 years ago
dear friend in triangle abc tana+tanb+tanc=tanatanbtanc you try this way in triangle a+b+c=180 then b+c=180-a apply tan on both sides
sreekanth sree
8 years ago
hello gowri sankar we know that the total angle in the triangle is 180,a+b+c=180,tan a+tan b+tan c=tan a tan b tan c.this is trignometric identities.
sudarshan
8 years ago
In triangle abc a+b+c=180 then tan a + tan b + tanc = tanatanbtanc.It is from trigonametric identity.
 
naresh raju
8 years ago
DEAR GOWRI,
IN the triangle A,B and C  A+B+C=180 then tanA+tanB+tanC=tanAtanBtanC.it could be solved by some trignometriv identities.
Dheeru chowdary
8 years ago
mr.gowri,
    it can be proved by using a+b+c=180.,
           and the answer is tanAtanBtanC which is an trignometry identity
 
Praveen
7 years ago
tan(a+b+c)=(tan a+tan b+tan c – tana*tanb*tanc) /(1 - tan a *tan b – tan*b tanc- tan a*tanc)
 since
     a+b+c=180, so tan(a+b+c)=0,
then
tan a+tan b+tan c – tana*tanb*tanc=0
tan a+tan b+tan c = tana*tanb*tanc